OpenStudy (anonymous):

The specific heat of silver is 0.057 calories/gram°C. If 10.0 grams of silver were heated and the temperature of the sample changed by 20.0°C, how many calories of heat energy were absorbed by the sample? 3,508 calories 0.029 calories 11.4 calories 0.114 calories

OpenStudy (aaronq):

have you seen this formula before? \(\sf \huge q=M*C_p*\Delta T\)

OpenStudy (anonymous):

ya i just don't under stand it

OpenStudy (aaronq):

can you define the terms in the equation?

OpenStudy (anonymous):

ya but just dont know how exactly how to solve it.

OpenStudy (aaronq):

you pretty much plug in the data from the question into the formula and solve. try to set the equation up and i'll come double-check it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0=10.0*0.057*20.0c

OpenStudy (aaronq):

q is what you're solving for, but the set up is perfect q=10.0*0.057*20.0

OpenStudy (anonymous):

q=11.4 calories